OpenStudy (anonymous):

a=21^3w, can someone help me solve for w?

OpenStudy (anonymous):

I have no idea how to solve, can someone show me the steps so I know how to solve those types of problems?

OpenStudy (anonymous):

take natural log on both sides then remove the power by getting back before the qoutient..

OpenStudy (anonymous):

Can you work it out I don't do well with words?

OpenStudy (anonymous):

it will be like this.. ln4=ln21^3w =>ln4=3wln21 =>ln4/3ln21=w =>w=ln4/ln21^3 =>w=ln(4-21^3)

OpenStudy (anonymous):

@keepsecure93

OpenStudy (anonymous):

In? what is In?

OpenStudy (anonymous):

You never included a into the equation. @zaibali.qasmi

OpenStudy (anonymous):

ln mean absolute log

OpenStudy (anonymous):

you didn't add in the "A" @zaibali.qasmi